[Brahms-l] Shift report 20050518 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Wed May 18 2005 - 08:09:07 EDT
Shifter leader was rk
Shifters were 


1. We had beam all night but for about 2 hours - the dump was unexpected and caused by failure of our D2 - 24V power supply failure and problems with waveform generator - will be fixed today.

2. The following runs were recorded (FS at 2.3deg full field B, MRS at 90deg full field B):
14733-35,14737-45 (14740-14742 non-existent because of IOserver and RunServer problems).

3. Once again - big problems with D2 (D1 is also not doing fine, as reported by previous shifter - HI) - few modules will be changed today. Probably D2 trips caused also DC to trip - had to restart 1458B mainframe.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed May 18 08:09:21 2005

This archive was generated by hypermail 2.1.8 : Wed May 18 2005 - 08:09:32 EDT